Define mean free path. Assuming that the
collision probability is a function of distance,
obtain the Survival Equation.
1
Expert's answer
2019-05-21T08:55:28-0400

The mean free path is the average distance travelled by a moving particle between successive collisions which modify its direction or energy or other particle properties.

The mean free path is equal to

"\lambda = \frac { 1}{\sqrt2 \pi d^2 n_V}"

Since each collision removes a particle from the group of uncollided particles, the change in Nc and dNc can be written as


"dN_C = -P_C(T, P)N_C dz (1)"

where PC(T, P) is constant defining that a collision occurs.

Integrating equation (1) using the boundary condition that when z=0, Nc=N, gives the fraction of particles at any position, z, that have yet to collide


"\frac { N_C}{N}=exp (-P_C(T, P) z) (2)"

The equation is called as survival equation it describes the number of particles that have survived travelling distance z without colliding.